Rewrite the quadratic equation in standard form by completing the square.
f(x)=x^2+8x+14
Answer (f(x)= (x, plus or minus ___)^2 plus or minus ____
I have worked this to x, 4 as well as x, -4. I have also come up with x, +8.
Also for the second answer I have some up with +14 or sqrt 2.

Found 2 solutions by nerdybill, scott8148:
Answer by nerdybill(7384)

Rewrite the quadratic equation in standard form by completing the square.
f(x)=x^2+8x+14
.
f(x)=x^2+8x+14
f(x)=(x^2+8x)+14
f(x)=(x^2+8x+16)+14-16
f(x)=(x+4)(x+4)-2
f(x)=(x+4)^2-2 (this is the standard form)

Answer by scott8148(6628)

x^2 + 8x + 16 is (x + 4)^2

when you add 2 to complete the square, you must also subtract 2 so the value of the expression stays constant

f(x) = x^2 + 8x + 14 + 2 - 2

f(x) = x^2 + 8x + 16 - 2

f(x) = (x + 4)^2 - 2 ___ this is called the vertex form of the equation, the vertex of the parabola is (-4,-2)